1.3.Blood glucose levels for obese patients have a mean of 100 with a standard deviation of 15. A
researcher thinks that a diet high in raw cornstarch will have a positive or negative effect on blood
glucose levels. A Sample of 30 patients who have tried the raw cornstarch diet has a mean glucose
level of 140.
Test the hypothesis that the raw cornstarch had an effect.
1
Expert's answer
2020-09-07T19:17:18-0400

Since the population standard deviation is known, Z-test is used.

"H_0:\mu=100"

"H_1:\mu\ne100"

Let "\alpha=0.05"

"Z=\frac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}"

"Z_{0.05}=1.96" (two tailed)

"Z=\frac{140-100}{\frac{15}{\sqrt{30}}}=14.61"

Since "Z=14.61>1.96" , we reject the null hypothesis and conclude that the raw cornstart had an effect at 5% level of significance.